2^30+2^30+2^30+2^30= ?
Answer by Boreal(15235)

2^30 is slightly more than a billion, so it is like adding 4 of them and getting 4 billion.
The actual answer is 4*(2^30)= (2^2)*(2^30)=2^32.
That is 4,294,967,296.